A room is seven ft longer than it is wide. If the perimeter is 46 ft, what are the dimensions of the room?
Answer by Cromlix(4381)

Width = x
Length = 7 + x
Perimeter = 2*length + 2*width
46 = 2(7 + x)+ 2x
46 = 14 + 2x + 2x
46 = 14 + 4x
4x = 46 - 14
4x = 32
x = 8
Width = 8 ft
Length = 15 ft
Hope this helps.
:-)